# Recent Advances in Discrete Speech Tokens: A Review

### v1

@@ III-A Model Architectures
Figure 4: Different neural architectures of acoustic tokens for reconstruction. Acoustic tokens are obtained after the quantization modules. Note that inputs and outputs can be waveforms, frequency-domain features or even SSL features depending on purpose and design.
@@ III-A1 VQ-GAN
VQ-GAN [89] is a very commonly adopted framework of acoustic tokens that trains a VQ-VAE with GAN objectives. Besides the original reconstruction and vector quantization objectives in a VQ-VAE, VQ-GAN uses discriminators $d_{\xi}(\bm{x},\hat{\bm{x}})$ to distinguish real and reconstructed data that adversarially trains the generator network of $f_{\theta},q_{\phi}$, and $g_{\psi}$. In acoustic tokens, there are usually multiple discriminators, e.g. multi-resolution and multi-scale STFT discriminators from the neural vocoder researches [90, 91]. The generator architecture of VQ-GAN-based acoustic tokens has multiple choices, with the three most representative onesvisualized in Fig.5: CNN-based, Transformer-based, and U-Net-based.
@@
The CNN-based generator is the most widely used architecture so far in acoustic tokens. SoundStream [61] and EnCodec [62] are two famous early neural acoustic tokens that operate in an end-to-end VQ-GAN manner. SoundStream is also the basis for Lyra V2 codec [^4] https://github.com/google/lyra. They receive time-domain waveforms as inputs and directly reconstruct waveforms. Their encoder and decoder have a mirrored architecture to perform down and up-samplings. In SoundStream, the encoder and decoder are purely constructed by convolutional neural networks (CNNs) while EnCodec augments them with an LSTM. The CNN encoder down-samples the waveform to a high-dimensional embedding sequence, whose frame rate is determined by the sampling rate, CNN kernel sizes and strides at a fixed ratio. The continuous embeddings are passed to an RVQ quantizer, and the quantized vectors are summed before being transformed to the waveform domain by the CNN decoder. The training criteria include reconstruction loss (in the time and frequency domain), adversarial loss, feature matching loss, and quantization losses for RVQ layers. To allow for a flexible choice of bitrates, structured dropout is adopted where the number of codebooks in the RVQ module can be randomly chosen [61], such that only a portion of quantizers in front are activated during training. The acoustic tokens can consequently reside in variable bitrates depending on the chosen number of RVQ quantizers. The inputs and outputs of the codec model can also be frequency-domain features like magnitude and phase spectra for reducing computation burden [67]. There, the convolution kernels are typically 2D instead of 1D in the time-domain codecs.
@@
Later, Transformers [3] have been adopted, e.g. in Single-Codec [71] and Mimi [10]. They can be directly applied to frequency-domain inputs and outputs. When operating on waveform-domain inputs or outputs, a CNN [10] or patchifying [19, 80] operation is usually added before or after the Transformer blocks. In Mimi, a shallow Transformer layer is added after the CNN-based encoder, and vice versa in its decoder. Recently, some propose to use purely Transformer-based backbone and discard the CNN blocks, e.g. TS3-Codec [19]. As Transformers are known for their superior modeling ability and scaling property, these works prove to outperform CNN-based codecs either with less computation [19] or larger scale [80]. However, to ensure stream-ability, an attention mask should be employed [10]. The encoder and decoder can also be designed to be different. For example, Single-Codec [71] uses Conformer [92] encoder and CNN decoder, while LSCodec [28] uses CNN encoder and Conformer decoder.
@@ III-B2 Approaches
Based on SoundStream and EnCodec, more codecs with advanced VQ methods, network structure, or optimization strategies have been researched with depth. HiFi-Codec [63] applies GRVQ on codecs to reduce the number of codebooks. SRCodec [69] proposes a dual attention mechanism and a split residual VQ strategy, which can also be regarded as GRVQ with interactions between groups. DAC [18] achieves remarkable reconstruction quality by adding periodic inductive bias, better discriminators, modified loss functions, and a better VQ mechanism from ViT-VQGAN [53]. Specifically, it performs L2-normed code lookup in a low-dimensional space (e.g. 8 or 32) instead of a high-dimensional space like 1024. Its VQ tricks are reported to improve code usage. APCodec [68] uses amplitude and phase spectra as inputs and outputs of the VQ-GAN codec model. Its encoder and decoder are based on an improved ConvNeXt v2 network [98], and it achieves fast and low-latency compression with a high sampling rate of 48kHz audio. Based on APCodec, [99] later reduces the necessary bitrate to 1kbps for high sampling rate scenarios by introducing additional bandwidth reduction and recovery modules before and after VQ-GAN. HILCodec [74] proposes spectrogram blocks, variance-constrained residual blocks and a multi-filter bank discriminator to achieve high fidelity and lightweight streaming codec. SuperCodec [75] improves the traditional CNN blocks in encoder and decoder into a selective down-sampling back-projection network for better performance. Mimi [10] also applies several training tricks to improve subjective perception: 1) introducing causal Transformers to both its encoder and decoder; 2) not applying VQ with a certain probability; 3) pure adversarial training without reconstruction losses.
@@
Additional training criteria around the VQ module are proposed for better utilization of the VQ codebook. ERVQ [100] proposes a more delicate code-vector replacement strategy and a codebook balancing loss to enhance the VQ usage. It also applies a similarity loss after consecutive RVQ layers to encourage each RVQ layer to focus on different speech features. Also, the training of neural speech codecs does not need to be end-to-end, i.e. the learning of VQ representations and signal reconstruction can be separated. AudioDec [66] adopts a two-stage training process that introduces adversarial losses and an additional vocoder after training only with metric losses. It achieves low latency while improving quality. APCodec+ [101] introduces this two-stage training process into APCodec, and declares that using adversarial loss throughout the entire process yields better performance.
@@
Although the aforementioned works all use regular GVQ or RVQ quantizers, other VQ methods also exist in speech codecs. NDVQ [79] improves the capacity of RVQ space by introducing stochasticity. NDVQ changes codebook vectors to isotropic Gaussian distributions parameterized by mean and variances. Instead of quantizing the continuous input to the closest codebook entry in each RVQ layer, NDVQ performs quantization by choosing the mean and variance with the greatest probability density. A sample is then drawn from the chosen Gaussian distribution as the VQ output, with a reparameterization technique. FSQ has also been introduced to speech codecs, like SQ-Codec [70] where the encoder constructs a low-dimensional space (32-dim), and each dimension is quantized by scalar rounding before being fed to the decoder. Stable-Codec [80] adopts FSQ in a Transformer-based architecture, exhibiting strong scalability to larger model sizes up to 950M parameter count. It also explores a flexible post-training quantization level adjustment technique and residual FSQ strategy. dMel [102] directly quantizes mel-filterbanks per dimension with evenly-paced boundaries between the minimum and maximum of filterbank values. This is similar to FSQ but is parameter-free.
@@
Note that previous acoustic tokens all require multiple quantizers, but single-codebook codecs have also been explored. Single-Codec [71] designs an encoder consisting of Conformer and bidirectional LSTM to compress mel spectrogram inputs. WavTokenizer [76] and BigCodec [77] further explores single-codebook codec modeling with better network designs or larger parameter count. TS3-Codec [19] adopts a fully Transformer design that leads to a better single-codebook codec with fewer computation overhead. LSCodec [28] also achieves single-codebook coding with speaker disentanglement (Section III-D). These single-codebook codecs with remarkably low bitrates offer great benefit to downstream speech generation models on simplicity and efficiency.
@@
Instead of capturing all the information through VQ layers like the previously mentioned codecs, some researchers have attempted to reduce the redundant bitrate of time-varying VQ codes. One straightforward method involves reducing the global information of an utterance in the time-varying codes by introducing a global encoder. The global information in speech, which includes speaker identity, channel effects, and other attributes, does not need to be repetitively encoded by time-varying discrete tokens. Disen-TF-Codec [20] is the first to explore VQ-GAN codec models that incorporate an additional encoder to output global features, thereby aiding the codec decoder in the reconstruction process. In Disen-TF-Codec, the global features are designed to be sequential to adapt to speaker changes during transmission. In TiCodec [21], the global tokens are time-invariant and vector-quantized instead. They are extracted from different segments of an utterance in conjunction with time-varying tokens. In CoFi-Codec [73], an ECAPA-TDNN [103] reference encoder is employed. The introduction of a global encoder also facilitates the development of single-codebook codecs, such as Single-Codec [71]. FreeCodec [81] further incorporates a prosody encoder [104] that compresses the low-frequency range of mel spectrograms into a low frame rate VQ sequence to assist in reconstruction.
Another typical example of temporal redundancy reduction is predictive coding, as seen in TF-Codec [56]. In this approach, the autoregressive prediction of latent features captures temporal-varying information in the latent space, which significantly reduces redundancy and entropy in the residual part for codebook quantization. LMCodec [105] employs autoregressive prediction from coarse codes (first RVQ levels) to fine codes (last RVQ levels) [5], enabling the transmission of fewer codes.
@@
Rather than relying solely on uni-resolution tokens, where all quantizers typically share the same temporal frequency of 25-86 Hz, it is reasonable to design multi-resolution codecs. This is because speech contains both fast and slow information streams. For instance, many vowels exhibit slowly changing characteristics, while events such as explosive consonants and background noises require fine-grained modeling. Therefore, incorporating multiple temporal resolutions in codecs is likely to reduce the necessary bitrate.
SNAC [22] is a notable multi-resolution acoustic token. It follows the DAC [18] architecture, but in each RVQ layer, residuals are downsampled before codebook look-up and upsampled afterward. For example, with a CNN downsampling factor of 512 on 24kHz waveforms, the first quantizer further downsamples the sequence by a factor of 4, and then upsamples the 12Hz quantized vectors by the same factor to compute quantization residuals. The downsampling factors for the second and third quantizers are 2 and 1, respectively. This design enables SNAC to outperform codecs with uni-resolution RVQ, especially under low bitrates. Similarly, LLM-Codec [24] adopts this method to achieve even lower frame rates with semantic distillation (Section III-C).
In contrast, CoFi-Codec [73] uses a U-Net architecture, where each encoder is a CNN with specific downsampling rates. The decoders follow a mirrored procedure, and quantizers are placed between each encoder-decoder pair with at a specific resolution. This configuration results in a multi-resolution representation, where, at each scale, GVQ is applied to the residual between the encoder and decoder hidden embeddings. ESC [72] adopts a similar approach where the encoders and decoders are constructed with Transformers. However, ESC modifies the frequency resolution in each layer rather than the time resolution.
In addition to multiple temporal resolutions, it is also feasible to consider the varying information intensities across different speech frames. Some frames carry critical information, while others may be less informative (e.g., silences). This observation motivates the design of codecs to allocate different numbers of quantizers for different speech frames. As an example, VRVQ [23] introduces a variable number of quantizers per frame in RVQ (i.e. the value $Q$) instead of using a fixed number. In VRVQ, a predictor receives the encoder’s hidden embeddings and outputs an importance map for each frame ranging from 0 to 1. This importance map determines the number of quantizers $Q$ for each frame and masks the quantizers beyond the first $Q$ quantizers. Since this masking process is not differentiable, surrogate functions are introduced to train the importance map predictor.
@@ III-C Acoustic Tokens with Semantic Distillation
Figure 6: Different semantic distillation methods in acoustic tokens.
@@ III-C1 Motivation
Acoustic tokens are a convenient choice for spoken language models, as they can be directly converted back to waveforms without the need for extra vocoders. However, if reconstruction is the sole objective of these tokens, their representation space may become overly complex and overly focused on acoustic details, in contrast to natural language tokens that primarily carry semantic information.. A natural improvement would be to incorporate speech semantic features either from speech self-supervised learning (SSL) models, supervised models, or even text transcriptions. Since speech SSL models aim to capture high-level phonetic or semantic information without external supervision [8], integrating SSL features does not impose additional data requirements for injecting semantic information into the training process. Acoustic tokens that incorporate criteria beyond reconstruction are sometimes referred to as having a “mixed objective” [1]. Given that the primary purpose of these models remains acoustic reconstruction in these models, we continue to refer to them as acoustic tokens. The process of introducing semantic information into acoustic tokens is termed semantic distillation, with approaches summarized in Fig. 6.
@@ III-C2 Approaches
Researchers have explored guiding some RVQ layers in acoustic tokens towards semantic features, which are typically SSL features. Since information in RVQ naturally follows a coarse-to-fine order, guiding early RVQ layers towards semantic-oriented features helps establish and reinforce a semantic-to-acoustic information hierarchy. For example, SpeechTokenizer [9] uses a HuBERT [32] SSL model to guide the first RVQ layer in EnCodec. This ensures that the first RVQ layer contains more semantic information, thereby pushing acoustic details to the subsequent RVQ layers. In SpeechTokenizer, this distillation is implemented either by regressing the first RVQ output to continuous HuBERT embeddings or by classifying it into discrete HuBERT tokens. The Mimi tokenizer [10] uses a WavLM [33] model as a semantic teacher and designs a specialized VQ module for distillation, rather than using the first RVQ layer. It claims to achieve a better semantic-acoustic trade-off compared to forcing acoustic information into the residual of the semantic quantizer. Since SSL feature guidance occurs only during the training stage, it does not incur additional inference costs. It has been reported that TTS language models trained with guided acoustic tokens exhibit better robustness than those on unguided acoustic tokens [9].
@@
A more direct approach to achieve semantic distillation is to integrate semantic knowledge into the codebook of quantizers. The encoder is then tasked with transforming the original speech into this semantic codebook space, while the decoder must learn to recover acoustics from this semantic space and the residuals. This method is explored in LLM-Codec [24] where pretrained automatic speech recognition (ASR) model Whisper [39], text language model T5 [106], and the LLM LLaMa-2 [107, 108] are employed as semantic teachers. LLM-Codec comprises three RVQ layers where all codebooks are initiated from the token embedding module of LLaMa-2 [108] and remain frozen during training. Specifically, the first RVQ codebook is constructed by selecting common words and average their corresponding sub-word embeddings from LLaMa-2. The rest two codebooks directly utilize the entire vocabulary space of LLaMa-2. Input vector sequences are downsampled at different rates before entering the first and second quantizers. The outputs of T5 and Whisper encoders are used to guide the first and second RVQ layers, respectively. This approach not only reduces the bitrate of the codec but also significantly enhances the semantic representation ability of LLM-Codec.
@@
Semantic features can also be compressed together with the speech waveforms or frequency features. This requires the encoder and quantizer to construct a shared acoustic and semantic space that balances the two information sources. The first attempt in this direction is made in [82] where Conformer representations from a pretrained wav2vec 2.0 [31] are combined with CNN encoder outputs for quantization. SemantiCodec [17], a codec designed for general audio, trains a quantizer to compress AudioMAE [109] features [^5] In fact, a stack of discretized and continuous AudioMAE features. without relying on acoustic inputs. The quantized SSL features then serve as a condition for acoustic reconstruction using latent diffusion, which resembles a vocoder that transforms semantic inputs into acoustic outputs. SoCodec [83] also directly quantizes semantic features from HuBERT, but incorporates a global acoustic condition to aid reconstruction, similar to TiCodec. With a downsampling semantic encoder, it remarkably explores a frame shift up to 240ms. Additionally, it has also been reported to reduce bitrate when aligned phoneme sequences are added to the encoder output before RVQ [67].
Moreover, semantic features can also serve as outputs, thereby reinforcing the constraint that semantic information be compressed into the discrete latent space. For instance, X-Codec [25] combines hidden HuBERT embeddings with acoustic features before RVQ and jointly optimizes acoustic and semantic reconstruction objectives. X-Codec 2.0 [84] improves it by using w2v-BERT 2.0 [110] and FSQ.

### v2

@@ III-A Model Architectures
Figure 4: Neural architectures of acoustic tokens. Note that inputs and outputs can be waveforms, frequency-domain features or even SSL features depending on purpose and design.
@@ III-A1 VQ-GAN
VQ-GAN [90] is a very commonly adopted framework of acoustic tokens that trains a VQ-VAE with GAN objectives. Besides the original reconstruction and VQ objectives in a VQ-VAE, VQ-GAN uses discriminators $d_{\xi}(\bm{x},\hat{\bm{x}})$ to distinguish real and reconstructed data that adversarially train the generator network composed of $f_{\theta},q_{\phi}$, and $g_{\psi}$. In acoustic tokens, there are usually multiple discriminators, e.g. multi-resolution and multi-scale STFT discriminators from the neural vocoder researches [91, 92]. The generator architecture of VQ-GAN-based acoustic tokens has multiple choices, with the three most representative ones visualized in Fig.5: CNN-based, Transformer-based, and U-Net-based.
@@
The CNN-based generator is the most widely used architecture so far in acoustic tokens. SoundStream [62] and EnCodec [63] are two famous early neural acoustic tokens that operate in an end-to-end VQ-GAN manner. They receive time-domain waveforms as inputs and directly reconstruct waveforms. Their encoder and decoder have a mirrored architecture to perform down and up-samplings. In SoundStream, the encoder and decoder are purely constructed by convolutional neural networks (CNNs) while EnCodec augments them with an LSTM. The CNN encoder down-samples the waveform to a high-dimensional embedding sequence, whose frame rate is determined by the sampling rate, CNN kernel sizes and strides at a fixed ratio. The continuous embeddings are passed to an RVQ quantizer, and the quantized vectors are summed before being transformed to the waveform domain by the CNN decoder. The training criteria include reconstruction loss (in the time and frequency domain), adversarial loss, feature matching loss, and quantization losses for RVQ layers. To allow for a flexible choice of bitrates, structured dropout is adopted where the number of codebooks in the RVQ module can be randomly chosen [62], such that only a portion of quantizers in front are activated during training. The acoustic tokens can consequently reside in variable bitrates depending on the chosen number of RVQ quantizers. The inputs and outputs of the codec model can also be frequency-domain features like magnitude and phase spectra for reducing computation burden [68]. There, the convolution kernels are typically 2D instead of 1D in the time-domain codecs.
@@
Later, Transformers [3] have been adopted, e.g. in Single-Codec [72] and Mimi [10]. They can be directly applied to frequency-domain inputs and outputs. When operating on waveform-domain inputs or outputs, a CNN [10] or patchifying [20, 81] operation is usually added before or after the Transformer blocks. In Mimi, a shallow Transformer layer is added after the CNN-based encoder, and vice versa in its decoder. Recently, some propose to use purely Transformer-based backbone and discard the CNN blocks, e.g. TS3-Codec [20]. As Transformers demonstrate superior modeling ability and scaling property, these works prove to outperform CNN-based codecs either with less computation [20] or larger scale [81]. However, to ensure stream-ability, an attention mask should be employed [10]. The encoder and decoder can also be designed to be different. For example, Single-Codec [72] uses Conformer [93] encoder and CNN decoder, while LSCodec [29] uses the reverse configuration.
@@ III-B2 Approaches
Based on SoundStream and EnCodec, more codecs with advanced VQ methods, network structure, or optimization strategies have been researched with depth. As an example, DAC [19] achieves remarkable reconstruction quality by adding periodic inductive bias, better discriminators, modified loss functions, and a better VQ mechanism from ViT-VQGAN [54] to improve codebook usage. Specifically, it performs L2-normed code lookup in a low-dimensional space (e.g. 8 or 32) instead of a high-dimensional space like 1024. Other architectural improvements include using frequency-domain inputs [69, 99, 72], variance-constrained residual blocks [75], multi-filter bank discriminator [75], selective down-sampling back-projection [76], etc.
@@
Several training tricks are explored, such as not applying VQ with a certain probability and pure adversarial training proposed in Moshi [10]. Also, the training of neural speech codecs does not need to be end-to-end, i.e. the learning of VQ representations and signal reconstruction can be separated. [67, 100] adopt a two-stage training process that introduces adversarial losses and an additional vocoder after training only with metric losses, to achieve improved quality. Additional training criteria around the VQ module are proposed for better VQ utilization, such as delicate code-vector replacement strategy, codebook balancing loss, and similarity loss between consecutive RVQ layers proposed in ERVQ [101].
@@
Other VQ methods besides GVQ or RVQ also exist in speech codecs. NDVQ [80] improves the capacity of RVQ space by changing codebook vectors to parameterized Gaussian distributions. FSQ has also been introduced to several speech codecs, like SQ-Codec [71] where scalar rounding is applied to each of its 32-dimensional latent space. Stable-Codec [81] adopts FSQ in a Transformer-based architecture, exhibiting strong scalability to large model sizes up to 950M parameter count. It also explores a flexible post-training quantization level adjustment technique and residual FSQ strategy.
@@
Note that most acoustic tokens require multiple quantizers, but single-codebook codecs have also been explored. Single-Codec [72] designs an encoder consisting of Conformer and bidirectional LSTM to better compress mel spectrogram inputs. WavTokenizer [77] and BigCodec [78] further explores single-codebook codec modeling with better network designs or larger parameter count. TS3-Codec [20] adopts a fully Transformer design that leads to a better single-codebook codec with fewer computation overhead. LSCodec [29] also achieves single-codebook coding with speaker disentanglement (Section III-D). These single-codebook codecs with remarkably low bitrates offer great benefit to downstream speech generation models on simplicity and efficiency.
@@
Instead of capturing all the information through VQ layers like the previously mentioned codecs, some researchers have attempted to reduce the redundant bitrate of time-varying VQ codes. One reasonable method is to encode the global information in speech, e.g. speaker timbre and channel effects, by a global encoder instead of the time-varying codes. Disen-TF-Codec [21] is the first to explore VQ-GAN codec models with an additional global encoder that aids the codec decoder. In Disen-TF-Codec, the global features are designed to be sequential to adapt to speaker changes during transmission. In TiCodec [22], the global tokens are time-invariant and vector-quantized instead. They are extracted from different segments of an utterance in conjunction with time-varying tokens. Similar global encoders are also seen in [84, 74, 72]. FreeCodec [82] further incorporates a prosody encoder [102] that compresses the low-frequency range of mel spectrograms into a low frame rate VQ sequence to assist in reconstruction.
Another typical example of temporal redundancy reduction is predictive coding, as seen in TF-Codec [57]. This approach captures temporal-varying information in the latent space by autoregressive prediction, which significantly reduces redundancy and entropy in the residual part for quantization. LMCodec [103] employs autoregressive prediction from coarse codes (first RVQ levels) to fine codes (last RVQ levels) [5], enabling the transmission of fewer codes.
@@
Rather than relying solely on uni-resolution tokens, where all quantizers share the same temporal frequency, it is reasonable to design multi-resolution codecs, because speech contains both fast and slow information streams. For instance, many vowels exhibit slowly changing characteristics, while events such as explosive consonants and background noises require fine-grained modeling. Therefore, incorporating multiple temporal resolutions in codecs is likely to reduce the necessary bitrate.
SNAC [23] is a notable multi-resolution acoustic token. It follows the DAC [19] architecture, but in each RVQ layer, residuals are downsampled before codebook look-up and upsampled afterward. This enables SNAC to have three RVQ streams at a frame rate of 12, 23, 47Hz respectively. Similarly, CoFi-Codec [74] achieves multi-resolution coding by GVQ quantizers within its U-Net-based architecture. LLM-Codec [25] also adopts this idea to achieve very low frame rates with semantic distillation (Section III-C).
In addition to multiple temporal resolutions, it is also feasible to consider the varying information intensities across different speech frames. This observation motivates the design of codecs to allocate different numbers of quantizers for different speech frames. As an example, VRVQ [24] automatically selects the number of RVQ quantizers per frame by a predictor that is jointly trained with the whole network.
@@ III-C Acoustic Tokens with Semantic Distillation
Figure 6: Different semantic distillation methods in acoustic tokens. Gray color indicates frozen during training.
@@ III-C1 Motivation
Acoustic tokens are a convenient choice for spoken language models, as they can be directly converted back to waveforms without the need for extra vocoders. However, if reconstruction is the sole objective of these tokens, their representation space may become overly complex and overly focused on acoustic details, in contrast to natural language tokens that primarily carry semantic information. A natural improvement is to incorporate speech semantic features either from speech self-supervised learning (SSL) models, supervised models, or even text transcriptions. Since speech SSL models aim to capture high-level phonetic or semantic information without external supervision [8], integrating SSL features does not impose additional data requirements for injecting semantic information into the training process. Acoustic tokens with criteria beyond reconstruction are sometimes referred to as having a “mixed objective” [1]. Given that the primary purpose of these models remains acoustic reconstruction in these models, we continue to refer to them as acoustic tokens. The process of introducing semantic information into acoustic tokens is termed semantic distillation, with approaches summarized in Fig. 6.
@@ III-C2 Approaches
The earliest effort in semantic distillation is to guide some RVQ layers in acoustic tokens towards semantic features, which are typically SSL features. Since information in RVQ naturally follows a coarse-to-fine order, guiding early RVQ layers towards semantic-oriented features helps establish and reinforce a semantic-to-acoustic information hierarchy. For example, SpeechTokenizer [9] uses a HuBERT [33] SSL model to guide the first RVQ layer in EnCodec. This ensures that the first RVQ layer contains more semantic information, thereby pushing acoustic details to the subsequent RVQ layers. This distillation is implemented either by regressing the first RVQ output to continuous HuBERT embeddings or by classifying it into discrete HuBERT tokens. LLM-Codec alternatively uses Whisper [40] and T5 [104] as semantic teachers. Mimi [10] uses a WavLM [34] teacher and applies distillation to a specialized VQ module rather than the first RVQ layer. Since SSL feature guidance occurs only during the training stage, it does not incur additional inference costs. It has been reported that TTS language models trained with such acoustic tokens exhibit better robustness than those with unguided tokens [9].
@@
A more direct approach to achieve semantic distillation is to integrate semantic knowledge into the codebook of quantizers. This forces the quantization space itself to be more semantic-related. This method is proposed in LLM-Codec [25] where all three RVQ codebooks are initiated from the token embedding module of LLaMa-2 [105] and remain frozen during training. This approach not only reduces the bitrate of the codec but also significantly enhances the semantic representation ability of LLM-Codec.
@@
Semantic features can also be compressed together with the acoustic features. This requires the encoder and quantizer to construct a shared acoustic and semantic space that balances the two information sources. The first attempt in this direction is made in [83] where Conformer representations from a pretrained wav2vec 2.0 [32] are combined with CNN encoder outputs for quantization. SemantiCodec [18] quantizes AudioMAE [106] SSL features without relying on acoustic inputs. The quantized SSL features then serve as a condition for acoustic reconstruction using latent diffusion, which resembles a vocoder that transforms semantic inputs into acoustic outputs. Providing aligned phoneme sequences instead of SSL features to the quantizer has also shown benefits on reducing bitrates [68].
Moreover, semantic features can also serve as outputs, thereby reinforcing the constraint that semantic information be compressed into the discrete latent space. For instance, [84, 26] combine hidden HuBERT embeddings with acoustic features before RVQ and jointly optimizes acoustic and semantic reconstruction objectives. X-Codec 2.0 [85] improves it by using w2v-BERT 2.0 [107] and FSQ.